A simple pendulum with a frequency of 0.056 Hz is as long as the largest known specimen of Pacific giant seaweed. What is this length?


1
Expert's answer
2021-11-04T10:15:09-0400

"T=2\pi\sqrt{l/g}\to l=gT^2/(4\pi^2)=9.81\cdot0.056^2/(4\cdot3.14^2)="


"=0.00078\ (m)=0.78\ (mm)"